let a1,a2,a3... be in A.P. and h1,h2,h3... be in H.P. . If a1=2=h1 and a30=25=h30 then a7h24+ a14+ a17=

Grade:11

1 Answers

Arun
6 years ago
for AP
a30 = a1 + 29 *d
25 = 2 +29d
d = 23/29
now we can find the value of a7 , a14 & a17.
 
Now if h1, h2, h3  are in HP then 1/h1 , 1/h2 , 1/h3 ….....will be in AP
now 
1/h30 = 1/h1 + 29* d1
by this we will find the value of d1.
and then we can find the value of h24.
 
now we have all the value which are required in question.
put the value and get the answer.
 
Hope it helps